pleas help me solve this question: A full grain silo empties at a constant rate, if the rate is 3.5 ft3/s. Write an expression to determine the amount of grain left after s seconds.
Answer by Theo(13342)

slope intercept form of a straight line equation is y = m * x + b.

m is the slope.
b is the y-intercept.

the slope is the same as he constant rate.

let y = the number of feet of grain left in the silo.

let x = the number of seconds.

you get y = -3.5 * x + b.

when x = s seconds, the equation becomes y = -3.5 * s + b.

b is the y-intercept which is the value of y when x = 0.

when x = 0, no seconds have elapsed and the silo is therefore at whatever height it was before you started to empty it.

that would be the value of b.

if we assume that the height of the grain in the silo was 150 feet, then the formula would become:

y = -3.5 * x + 150

after s seconds, the formula would become y = -3.5 * s + 150.

if s was equal to 20, then the formula would become y = -3.5 * 20 + 150 which would then become y = -70 + 150 which would then become y = 80.

after 20 seconds, there would be 80 feet of grain left in the silo.

the graph of the equation would look like this.



from the graph, you can see that the height of the grain in the silo is 150 feet when you start and gets down to 80 feet after 20 seconds and that the silo will be empty somewhere between 42 and 43 seconds.